ADDENDUM TO ASYMPTOTIC STABILITY IN DIFFERENTIAL EQUATIONS WITH UNBOUNDED DELAY

T.A. Burton  Alfredo Somolinos
Northwest Research Institute  Department of Mathematics
732 Caroline Street  Mercy College
Port Angeles, WA 98362  Dobbs Ferry, New York 10522

This addendum concerns the paper of the above title found in EJQTDE/1999/199913. Throughout that paper was the tacit assumption that the coefficient functions h(t), b(t), and C(t) are all continuous on their respective domains. Every result, as well as the existence result stated at the end of the first section, depended on those functions being continuous. A search of the paper indicates that we failed to state this. We regret any inconvenience which this may have caused any reader.

EJQTDE, 2001 No. 3, p. 1

1991 Mathematics Subject Classification. 34K20.

Key words and phrases. Liapunov Stability, Functional Equations, Unbounded Delay, Volterra Equations.